# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v1

@@ Abstract
Fine-tuning large language models (LLMs) using diverse datasets is crucial for enhancing their overall performance across various domains. In practical scenarios, existing methods based on modeling the mixture proportions of data composition often struggle with data whose domain labels are missing, imprecise or non-normalized, while methods based on data selection usually encounter difficulties in balancing multi-domain performance. To address these challenges, in this paper, we study the role of data diversity in enhancing the overall abilities of LLMs by empirically constructing contrastive data pools and theoretically deriving explanations for both inter- and intra-diversity. Building upon the insights gained, we propose a new method that gives the LLM a dual identity: an output model to cognitively probe and select data based on diversity reward, as well as an input model to be tuned with the selected data. Extensive experiments show that the proposed method notably boosts performance across domain-undetermined data and a series of foundational downstream tasks when applied to various advanced LLMs. We release our code and hope this study can shed light on the understanding of data diversity and advance feedback-driven data-model co-development for LLMs.
Machine Learning, ICML
@@ 1 Introduction
The advancement of large language models (LLMs), exemplified by open-source models such as the Llama3 series (Dubey et al., 2024), Qwen2 series (Yang et al., 2024), and the DeepSeek-V3 series (Liu et al., 2024a), along with closed-source models such as GPT-4 (Achiam et al., 2023), has revolutionized artificial intelligence by enhancing capabilities in foundational abilities such as common sense, reasoning, mathematics, and coding. Fine-tuning these models further optimizes their usability by enhancing their performance and aligning with specific human instructions and preferences  (Taori et al., 2023; Rafailov et al., 2023).
@@
To cultivate comprehensive capabilities in LLMs, fruitful studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  (Li et al., 2024b; Chen et al., 2024b; Qin et al., 2024; Zhao et al., 2024a). For example, methods focusing on data selection (Li et al., 2024a; Xia et al., 2024; Wang et al., 2023) and mixture (Ge et al., 2024) demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  (Lu et al., 2024; Liu et al., 2024b).
@@
However, real-world applications frequently encounter unlabeled data and difficulties in domain labeling (Ge et al., 2023), posing challenges for data mixture methodologies that take the domain tags as a prior, as well as the data selection approaches, which often prioritize quality over diversity especially with data sourced from quite different domains.
To leverage the best of both worlds, in this work, we propose a new fine-tuning method for LLMs named DaaR, which encourages the given model to learn diversity as a reward signal, and then to autonomously select domain-undetermined training data to achieve a theoretically informed trade-off to maximize diversity for overall model performance enhancement.
Our investigation begins by diving into the semantic diversity of LLM data, considering how to explicitly model it and how it can influence the overall performance of LLMs. We systematically construct contrastive data pools and conduct extensive empirical examinations of different distributional performances across various foundational LLM capabilities. Based on these observations, we provide theoretical discussions on a general assumption that a mixture of underlying component LLMs determines the data mixing effect, and the explanations regarding inter- and intra-diversity modeled from a data semantic perspective.
Figure 1: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B. (a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
Our theoretical insights illuminate that optimal solutions for overall model performance enhancement cannot be explicitly applied to domain-undetermined data. To address this, we integrate an external multi-layer perceptron (MLP) structure into the LLM to be tuned, acting as a trainable probe module conditioned on the inherent knowledge and weights of the given LLM. The probe module takes semantic entropy as a proxy measure of diversity, and output reward scores, and select suitable data subsets aligned with the LLM’s underlying mixture distributions for better diversity. This is achieved by training the probe module on synthetically generated domain-aware data by the model itself, and using the data after selection to further self-fine-tuning. Extensive experiments on various state-of-the-art (SOTA) LLMs like the Qwen and LLaMA series verify the effectiveness of DaaR in enhancing overall model capabilities over many SOTA baseline methods.
Our contributions are threefold:
  • We empirically and theoretically demonstrate how data plays a crucial role to enhance LLMs in semantic diversity, providing generalized modeling based on mixture of component models.
  • We propose a lightweight, self-rewarding method for selecting data in domain-undetermined scenarios, which utilizes synthesized model-aware data to maximize domain diversity and fine-tuning performance.
  • We verify that the proposed methods can effectively balance seven diverse benchmarks for improved overall performance when applied to representative LLMs, whereas other SOTA methods struggle in such challenging scenarios. Our code is open-sourced  [^1] https://github.com/modelscope/data-juicer/tree/DaaR to foster more in-depth understanding and future progress.
2 Related Work
@@
Fine-tuning is a pivotal training paradigm for enhancing LLMs’ task-specific and domain-specific capabilities. Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively (Zhou et al., 2024; Chen et al., 2024a). This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm (Paul et al., 2021) and token length (Raffel et al., 2020). More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty (Li et al., 2024a), GPT scoring (Chen et al., 2024d; Wettig et al., 2024), data model selection (Engstrom, 2024), and influence scores derived from loss (Xia et al., 2024).
@@
A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning (Yu et al., 2022; Ding et al., 2023). Data diversity encompasses aspects such as data deduplication (Abbas et al., 2023), the coverage scope of tags (Lu et al., 2024), model-based diversity evaluation (Liu et al., 2024b; Zhang et al., 2024), and scaling properties (Song et al., 2024). Typically, data mixing approaches (Albalak et al., 2023; Ye et al., 2024; Ge et al., 2024) focus on adjusting the proportional weights of different domains to enhance model capabilities.
@@
From a modeling perspective, our work shares a focus with many data mixing methodologies regarding diversity measurement. However, we address a more challenging setting relevant to real-world applications, where available datasets often lack domain labels. This situation presents challenges for existing mixing methods in balancing label granularity and normalization, and in determining which domains should be considered as candidates for improved solution spaces and tuning performance.
Methodologically, our approach aligns with the data selection category but differs in both the learning process (utilizing a rewarding model) and the selection criteria (employing semantic entropy). Additionally, we revisit the applicability of existing selection methods in multi-domain scenarios. This area is under-explored and requires new insights to avoid biased modeling of diverse domain distributions—a topic we further investigate in this paper.
3 Motivational Observations
In this section, we will explore how data with different domain-specific diversity affects model capabilities through experimental data pools, featuring contrastive data distributions in terms of inter-domain and intra-domain diversity.
@@ 3.1 Seed Data Pools and Basic Setting
Data Pools and Sources
To explore how selecting data samples from extensive and diverse data repositories affects the foundational capabilities of LLMs, we construct various data pools and consistently fine-tune LLMs on them. We then analyze the performance changes and attribute these changes to the different controlled data pools.
The seed data pool is sourced from following datasets: Dolly-15k (Conover et al., 2023) for common sense, Cot-en (Chung et al., 2024) for reasoning, Math-Instruct (Yue et al., 2024) for mathematics, and Code-Alpaca (Chaudhary, 2023) for coding. Each dataset was randomly tailored to 10,000 entries, resulting in a combined data pool of 40,000 entries. Following instruction tuning practices (Zhou et al., 2024; Liu et al., 2024b), we then uniformly sample 8,000 data entries as a referenced data pool for random baseline. In subsequent sections, we will introduce how to construct other comparative pools with the same size to random pool.
Benchmarks
Aligned with representative capabilities of leading open-source LLMs, we select the following widely used evaluation sets: NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) for common sense, Hellaswag (Zellers et al., 2019) for reasoning, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) for mathematics, MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) for coding. To evaluate the comprehensive performance of LLMs across domains, we employ the average metric (Avg) as the primary evaluation criterion.
Models & Implementation
To ensure the effectiveness and applicability of our empirical findings, we employ the Qwen2 series (Qwen2-7B & Qwen2.5-7B) (Yang et al., 2024) and the Llama3.1-8B (Dubey et al., 2024) as representative SOTA base models to be fine-tuned. All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix B.
@@ 3.2 Data Pools with Contrastive Distributions
Let the composite dataset $\mathcal{D}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$ comprise $K=4$ distinct domains, where each domain subset $\mathcal{D}_{k}$ contains $N_{k}=|\mathcal{D}_{k}|$ samples. We represent each data instance through its semantic embedding $\mathbf{x}_{i}^{(k)}\in\mathbb{R}^{d}$ extracted from the Embedding layer of the pretrained LLM, capturing high-dimensional semantic features. The domain centroid $\mathcal{C}_{k}$ serves as the semantic prototype:
$\small\mathcal{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}^{(k)}.$ (1)
This centroid-based representation enables geometric interpretation of domain characteristics in the embedding space. We dissect data diversity into two complementary aspects:
@@
It quantifies the diversity between distinct domains through centroid geometry. For sample $\mathbf{x}_{i}^{(k)}$, its cross-domain similarity is measured by:
$\small\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})=\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{j}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{j}\|}.$ (2)
The global inter-diversity metric $\Phi_{\text{inter}}$ computes the expected pairwise centroid distance:
$\small\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right]=\frac{1}{\binom{K}{2}}\sum_{k=1}^{K-1}\sum_{l=k+1}^{K}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}.$ (3)
This formulation reflects a key insight: maximizing $\Phi_{\text{inter}}$ encourages domain separation, while minimization leads to overlapping representations. Fig. 1 demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig. 1.(c)), whereas low values produce entangled distributions (Fig. 1.(d)). Full analysis is detailed in Appendix D.1.
@@
Focusing solely on the separation between different domains may hinder the model’s ability to learn the knowledge specific to a given domain. Hence we measure variation within each domain. We calculate sample similarity to its domain center:
$\small\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})=\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{k}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{k}\|}.$ (4)
And the domain-level variance metric is defined as:
$\small\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (5)
Controlled manipulation of $\Phi_{\text{intra}}$ reveals critical trade-offs: lower variance (tight clusters near $\mathcal{C}_{k}$) enhances domain-specific feature learning but risks over-specialization. Higher variance improves robustness at the cost of potential cross-domain interference. The visualization in Fig. 1 (e-f) illustrates this scenario that concentrated distributions exhibit sharp marginal peaks, while dispersed variants show overlapping density regions.
@@
The 20% interval results in five choices of data selection per domain, parameterizing the trade-off between diversity preservation and domain specificity. As demonstrated in Appendix D.2 (Fig. 10 and Fig. 11), this quantization process induces measurable distribution shifts.

### v2

@@ Abstract
Fine-tuning large language models (LLMs) using diverse datasets is crucial for enhancing their overall performance across various domains. In practical scenarios, existing methods based on modeling the mixture proportions of data composition often struggle with data whose domain labels are missing, imprecise or non-normalized, while methods based on data selection usually encounter difficulties in balancing multi-domain performance. To address these challenges, in this work, we investigate the role of data diversity in enhancing the overall abilities of LLMs by empirically constructing contrastive data pools and theoretically deriving explanations. Building upon the insights gained, we propose a new method that gives the LLM a dual identity: an output model to cognitively probe and select data based on diversity reward, as well as an input model to be tuned with the selected data. Extensive experiments show that the proposed method notably boosts performance across domain-undetermined data and a series of foundational downstream tasks when applied to various advanced LLMs. We release our code and hope this study can shed light on the understanding of data diversity and advance feedback-driven data-model co-design for LLMs.
11footnotetext: Equal contribution.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) — exemplified by open-source series such as LLaMA llama3series , Qwen qwen2series , and DeepSeek deepseekv3 , as well as closed-source models like GPT gpt4 — has transformed artificial intelligence by significantly enhancing capabilities in core areas such as common sense reasoning, mathematics, and code generation. Fine-tuning these models further improves their usability by aligning their behaviors with specific instructions and human preferences alpaca2023 ; dpo .
@@
To cultivate comprehensive capabilities in LLMs, prior studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  li2024quantity ; djv2 ; qin2024surveycodev ; zhao2024beyond . For example, methods focusing on data selection li2024super ; xialess ; wang2023far and mixture ge2024bimix demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  lu2023instag ; liu2023deita .
@@
However, real-world applications frequently encounter unlabeled data and difficulties in domain labeling ge2023openagi , posing challenges for data mixture methodologies that take the domain tags as a prior, as well as the data selection approaches, which often prioritize quality over diversity, especially with data sourced from quite different domains.
To leverage the best of both worlds, in this work, we dive into the role of semantic diversity of LLM data. Our investigation begins with a systematic analysis of 40k instruction pairs across four capability domains. Through controlled experiments with contrastive data pools, we uncover two insights: (1) Optimal diversity thresholds for model performance vary significantly across architectures and task domains; (2) Current centroid-based diversity metrics lu2023instag fail to capture the dynamic interaction between inter-domain separation and intra-domain variance when labels are unavailable.
These findings motivate our theoretical analysis, which establishes that traditional data mixing strategies assuming known domain weights ye2024data cannot achieve Pareto-optimal performance in label-free settings. We prove that when domain labels are unavailable, the optimal data selection strategy must simultaneously maximize the mutual information between selected samples and preserve the base model’s latent feature geometry.
We operationalize these insights through DaaR, a self-supervised framework that learns diversity as a reward signal through three technical innovations: (1) Automatic synthesis of model-aware domain centroids via iterative embedding-space generation, (2) A lightweight MLP probe trained to predict semantic entropy using only the LLM’s frozen embeddings, and (3) Closed-loop fine-tuning where the model’s own diversity estimates guide subsequent data selection.
Extensive evaluations across 7 benchmarks and 3 model families reveal that DaaR achieves new state-of-the-art (SOTA) average performance, consistently outperforming 9 baseline methods on high-difficulty tasks while maintaining computational efficiency. Crucially, our method requires no domain labels or external models - the LLM self-supervises its diversity estimation through geometric constraints in its native embedding space.
Our core contributions for LLM data optimization are three-fold:
  • Formal characterization of diversity’s dual role in LLM fine-tuning through controlled experiments with contrastive data pools.
  • A new method with label-free diversity reward mechanism using embedding-space entropy prediction, theoretically grounded in mutual information maximization.
  • Extensive empirical validations showing consistent gains across model families and tasks, with notable improvements in mathematical reasoning (+27%) and coding (+7.4%), whereas other baselines struggle in such challenging scenarios. Our code is released at https://github.com/modelscope/data-juicer/tree/DaaR to foster more data-centric research for LLMs.
2 Preliminaries
2.1 Related Works
@@
Fine-tuning is a pivotal training paradigm for enhancing LLMs’ domain-specific capabilities. Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively zhou2024lima ; djv1 . This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm paul2021el2n and token length raffel2020exploring . More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty li2024super , GPT scoring chen2024alpagasus ; 2024qurating , data model selection 2024dsdm , and influence scores derived from loss xialess .
@@
A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning yu2022can ; ding2023enhancing . Data diversity encompasses aspects such as data deduplication abbas2023semdedup , the coverage scope of tags lu2023instag , model-based diversity evaluation liu2023deita ; zhang2024harnessing , and scaling properties song2024scaling . Typically, data mixing approaches albalak2023efficient ; ye2024data ; ge2024bimix focus on adjusting the proportional weights of different domains to enhance model capabilities.
@@
This work relates to data mixing and selection fields through its focus on diversity quantification, but advances both of them in three key aspects: First, we address practical challenges in real-world applications where unlabeled datasets hinder existing methods’ ability. Second, our framework distinguishes by integrating reward-driven learning with semantic entropy-based selection criteria. Third, we critically examine the under-explored applicability of current selection paradigms in multi-domain contexts, revealing critical limitations in modeling diverse distribution patterns.
2.2 Problem Setup
In this paper, we consider learning from domain-undetermined data mixtures to enhance LLM downstream performance in target domains. Let $\{\mathcal{D}_{k}\}_{k=1}^{K}$ denote $K$ target domain distributions, where each $\mathcal{D}_{k}:\mathcal{X}\rightarrow[0,1]$ defines a probability distribution over samples $x\in\mathcal{X}$. For model parameters $\theta$, the domain-specific performance is measured by $\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta)]$. Our objective is to minimize the weighted multi-domain loss: $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta)],$ where $\lambda_{1:K}$ are non-negative weights with $\sum_{k=1}^{K}\lambda_{k}=1$, reflecting application-specific priorities.
For each domain $\mathcal{D}_{k}$, we assume only a natural language description $d_{k}$ is available while NO target domain samples are accessible, contrasting with conventional data selection methods requiring labeled examples xialess ; xie2023data ; renduchintala2024smart ; ye2024data in LLM training. We are provided with an domain-unlabeled dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ containing domain-agnostic samples. A straightforward approach is direct fine-tuning using all $\mathcal{D}$, but it often yields suboptimal results (empirically evidenced in Table 2). Instead, we propose selecting a subset via binary vector $\alpha\in\{0,1\}^{N}$, where $\alpha_{i}=1$ indicates $x_{i}$ is selected. The LLM is trained by solving: $\min_{\theta}\sum_{i=1}^{N}\alpha_{i}\ell(x_{i};\theta)$, leading to the bi-level optimization:
$\small\min_{\alpha}\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{x\sim\mathcal{D}_{k}}[\ell(x;\theta^{*}(\alpha))],\quad\text{s.t.}\ \ \theta^{*}(\alpha)=\arg\min_{\theta}\sum_{i=1}^{N}\alpha_{i}\ell(x_{i};\theta).$ (1)
3 Diversity: A Critical Factor for LLM Fine-Tuning
In this section, we investigate how domain-specific diversity in labeled training data modulates model capabilities through empirical analysis of well-designed data pools, complemented by a theoretical perspective that explains the observation and provides mechanistic insights.
Figure 1: The illustration highlights our observations and the proposed DaaR method. Observation shows the t-SNE visualization of embeddings for data samples with different distributions, leading to varying LLM’s evaluation performance. In the domain-undetermined data selection scenario, DaaR introduces a dual-identity framework for LLMs: ① An output model to probe and select data based on diversity reward, ② An input model to be tuned with the selected data. The diversity probe is trained on model-aware synthetic data, enabling domain discrimination and diversity reward prediction.
@@ 3.1 Seed Data Pools and Basic Setting
Data Pools and Sources To explore how selecting data samples from extensive and diverse data repositories affects the foundational capabilities of LLMs, we construct various data pools and consistently fine-tune LLMs on them. The seed data pool is sourced from the following datasets: Dolly-15k dolly-15k for common sense, Cot-en cot-en for reasoning, Math-Instruct math-instruct for mathematics, and Code-Alpaca code-alpaca for coding. Each dataset was randomly tailored to 10,000 entries, resulting in a combined data pool of 40,000 entries. Following instruction tuning practices zhou2024lima ; liu2023deita , we then uniformly sample 8,000 data entries as a reference data pool for the random baseline.
Benchmarks Aligned with representative capabilities of leading open-source LLMs, we select the following widely used evaluation sets: NQ nq and TriviaQA triviaqa for common sense, Hellaswag hellaswag for reasoning, GSM8K gsm8k and MATH math for mathematics, MBPP mbpp and HumanEval humaneval for coding. To evaluate the comprehensive performance of LLMs across domains, we employ the average metric (Avg) as the primary evaluation criterion.
Models & Implementation We employ the Qwen2 series (Qwen2-7B & Qwen2.5-7B) qwen2series and the Llama3.1-8B llama3series as representative SOTA base models to be fine-tuned.
All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix B.
@@ 3.2 Data Pools with Contrastive Distributions
Each domain dataset $\mathcal{D}_{k}$ contains $N_{k}=|\mathcal{D}_{k}|$ samples, and we represent each data sample through its semantic embedding $\mathbf{x}_{i}^{(k)}\in\mathbb{R}^{d}$ extracted from the embedding-layer of the pretrained LLM. The domain centroid $\mathcal{C}_{k}$ serves as the semantic prototype $\mathcal{C}_{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbf{x}_{i}^{(k)}$. This centroid-based representation enables geometric interpretation of domain characteristics in the embedding space. We dissect data diversity into two complementary aspects:
@@
It quantifies the diversity between distinct domains through centroid geometry. For sample $\mathbf{x}_{i}^{(k)}$, its cross-domain similarity is measured by $\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})$. The global inter-diversity metric $\Phi_{\text{inter}}$ computes the expected pairwise centroid distance:
$\phi_{\text{inter}}(\mathbf{x}_{i}^{(k)})=\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{j}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{j}\|},\quad\Phi_{\text{inter}}=\mathbb{E}_{k\neq l}\left[\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}\right]=\frac{1}{\binom{K}{2}}\sum_{k=1}^{K-1}\sum_{l=k+1}^{K}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|_{2}.$ (2)
This formulation reflects a key insight: maximizing $\Phi_{\text{inter}}$ encourages domain separation, while minimization leads to overlapping representations. Observation of the upper part of Fig. 1 (full-size version is in Appendix, Fig. 16) demonstrates this continuum through t-SNE projections – high $\Phi_{\text{inter}}$ manifests as distinct cluster separation with clear margins (Fig. 16.(c)), whereas low values produce entangled distributions (Fig. 16.(d)). Full analysis is detailed in Appendix H.1.
@@
Focusing solely on the separation between different domains may hinder the model’s ability to learn the knowledge specific to a given domain. Hence we calculate sample similarity to its domain center in $\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})$, and the domain-level variance metric is defined as $\Phi_{\text{intra}}^{(k)}$.
$\phi_{\text{intra}}(\mathbf{x}_{i}^{(k)})=\frac{\mathbf{x}_{i}^{(k)}\cdot\mathcal{C}_{k}}{\|\mathbf{x}_{i}^{(k)}\|\|\mathcal{C}_{k}\|},\quad\Phi_{\text{intra}}^{(k)}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\|\mathbf{x}_{i}^{(k)}-\mathcal{C}_{k}\|_{2}^{2}.$ (3)
Controlled manipulation of $\Phi_{\text{intra}}$ reveals critical trade-offs: lower variance enhances domain-specific learning but risks over-specialization, while higher variance improves robustness with the risk of cross-domain interference. The visualization in Fig. 16 (e-f) illustrates this scenario that concentrated distributions exhibit sharp marginal peaks, while dispersed variants show overlapping density regions.
@@
The 20% interval results in five choices of data selection per domain, parameterizing the trade-off between diversity preservation and domain specificity. As demonstrated in Appendix H.2 (Fig. 18 and Fig. 19), this quantization process induces measurable distribution shifts.